1.   A flux density of 1.5T is produced in a piece of cast steel by a magnetizing force of 1250A/m. Find the relative permeability of the steel under these conditions.


2.   Determine the magnetic field strength required to produce a flux density of 0.3T in an air gap of length 15 mm. 



1
Expert's answer
2021-11-24T12:00:08-0500

1.

"B = \mu_0 \mu_γH \\

\mu_γ = \frac{B}{\mu_0 H} \\

= \frac{1.5}{4 \pi \times 10^{-7} \times 1250} \\

= 955.4"

2. B = 0.30 T

"B = \mu_0H \\

\mu_0 = 4 \pi \times 10^{-7} \\

H = \frac{0.3}{4 \pi \times 10^{-7}} \\

H = 199044 \; A/m"

Magnetomotive force "(MMF) = H \times L"

L = 15 mm "= 15 \times 10^{-3} \;m"

"MMF = 199044 \times 15 \times 10^{-3} = 2985.6 \; AT"